OpenStudy (anonymous):

Franco bought 1.5 pounds of deli cheese for six dollars and seventy-five cents. If he went back for another 2 pounds of the same cheese, how much could Franco expect to pay? (Now, my question is, how do I set this up into a proportion. I don't need to know the answer to the word problem above, I just need to know how I can set it up into a proportion. Once when I can do that then I'm good for solving it on my own.)

OpenStudy (triciaal):

1.5 lbs cost 6.75 what is the cost for 2 lbs? let the cost = x option 1 1.5 * 1lb cost 6.75 divide by 1.5 1 lb cost 6.75/1.5= 2 lbs cost 2* 1 lb = option 2 1.5/6.75 = 2/x solve for x = 2 * 6.75/1.5 =

OpenStudy (anonymous):

Oh, I get it now. Thanks, that's actually all I needed. I got from here. :)